OpenStudy (anonymous):

Allie is running around a track at a rate of about 4 meters per second. What is her rate in feet per minute

OpenStudy (anonymous):

1 inch = 2.54 centimeter

OpenStudy (anonymous):

240/m = 4 * 60(seconds) = 240m per minute

OpenStudy (anonymous):

Allie is running around a track at a rate of about 4 meters per second. What is her rate in feet per minute? A. about 18.3 feet per minute B. about 73.2 feet per minute C. about 787 feet per minute D. about 3149 feet per minute

OpenStudy (anonymous):

those are the answers

OpenStudy (anonymous):

1 foot = 12 inch = 12 * 2.54 cm = ( 12 * 2.54 / 100 ) meter

OpenStudy (anonymous):

60 seconds = 1 minute

OpenStudy (anonymous):

C is the answer

OpenStudy (anonymous):

240 meters converted to feet = 787.402 feet

OpenStudy (anonymous):

I drew up a visual so you can see how it works. Just plug in the questions variables in the first spot, then look up the conversions for the second and third and do the multiplication.

OpenStudy (anonymous):

Thanks for the medal!

OpenStudy (anonymous):

yea